# RuleRAG: Rule-Guided Retrieval-Augmented Generation with Language Models for Question Answering

### v1

@@ Abstract
Retrieval-augmented generation (RAG) footnotetext: * Corresponding Authors framework has shown promising potential in knowledge-intensive question answering (QA) by retrieving external corpus and generating based on augmented context. However, existing approaches only consider the query itself, neither specifying the retrieval preferences for the retrievers nor informing the generators of how to refer to the retrieved documents for the answers, which poses a significant challenge to the QA performance. To address these issues, we propose Rule-Guided Retrieval-Augmented Generation with LMs, which explicitly introduces symbolic rules as demonstrations for in-context learning (RuleRAG-ICL) to guide retrievers to retrieve logically related documents in the directions of rules and uniformly guide generators to generate answers attributed by the guidance of the same set of rules. Moreover, the combination of queries and rules can be further used as supervised fine-tuning data to update retrievers and generators (RuleRAG-FT) to achieve better rule-based instruction following capability, leading to retrieve more supportive results and generate more acceptable answers. To emphasize the attribution of rules, we construct five rule-aware QA benchmarks, including three temporal and two static scenarios, and equip RuleRAG with several kinds of retrievers and generators. Experiments demonstrate that training-free RuleRAG-ICL effectively improves the retrieval quality of +89.2% in Recall@10 scores and generation accuracy of +103.1% in exact match scores over standard RAG on average across the five benchmarks, and further fine-tuned RuleRAG-FT consistently yields more significant performance enhancement. Extensive analyses indicate that RuleRAG scales well with increasing numbers of retrieved documents and exhibits generalization ability for untrained rules. Our code and benchmarks are available at https://github.com/chenzhongwu20/RuleRAG_ICL_FT.
@@ 1 Introduction
Large language models (LLMs) have achieved impressive language generation capability and excelled as knowledge learners for their well-known in-context learning (ICL) ability (Brown et al., 2020; Ouyang et al., 2022; Chowdhery et al., 2024). Despite the success, the full-parametric knowledge stored in LLMs requires substantial computational costs to keep their memory up-to-date and struggles to precisely manipulate fine-grained queries, especially in knowledge-intensive tasks (Jiang et al., 2023c; Shao et al., 2023). As complementary, RAG represents a novel framework that integrates LLMs with non-parametric information and injects the retrieved knowledge in a plug-and-play manner (Lewis et al., 2020; Dhingra et al., 2022). By explicitly decoupling the knowledge retrieval phase from the answer generation phase, RAG exhibits superior performance in many NLP tasks, such as open-domain QA (Trivedi et al., 2023) and natural language inference (Qin et al., 2023).
@@
Figure 1: (a) Without the help of rules, the current RAG can only retrieve relevant documents at the shallow semantic level, rather than the overall semantics of the query, and thus get confused in answering. (b) Guided by the rule $r$ related to the query, our proposed RuleRAG first retrieves supportive documents that are logically related to the query and then attributes the correct answer, “France”.
@@
However, two high-level issues exist in the current RAG frameworks. First, in the retrieval phase, the imperfect retrieval component can not guarantee that the recalled information will always be the most pertinent and helpful to the queries. The reason is that the retrievers in retrieval-augmented language models (RALMs) are mostly trained on unsupervised text (Izacard et al., 2024) or trained end-to-end (Guu et al., 2020; Borgeaud et al., 2022a), leading to their insufficiency in retrieving the necessary statements for reasoning (BehnamGhader et al., 2023). Secondly, in the generation phase, the LLMs in the current RAG are not specifically informed of how to exploit noisy retrieved content properly, since relationships between a wide range of facts are rarely explicitly “pointed out” and “supervised” in the pre-training corpora of LLMs. For example, REPLUG (Shi et al., 2023) and In-Context RALM (Ram et al., 2023) fuse off-the-shelf LLMs with generic retrievers and treat LLMs as black boxes. Even if answered correctly, they still lead to implicit attribution processes that are difficult to explain and verify. Therefore, these RAG frameworks are neither inherently trained to retrieve along reasonable retrieval directions nor organically attribute retrieved content to answers.
In contrast to these existing RAG frameworks, we observe that widespread logical rules can explicitly guide people to accomplish a given task. For example, for the addition of large numbers, a math problem, humans can easily solve the addition of any two numbers after learning the rules of column addition. To implement rule-based calculating,  Hu et al. (2024) instructs transformers to cite basic addition rules by explicitly fine-tuning transformers with them. While answering factual queries by means of retrieve-then-read in our task, human experts spontaneously search for relevant information as intermediate results following a priori rules before answering and refer to these rules again while deciding final answers (Shnarch et al., 2020). As shown in Figure 1 (a), if we ask $\mathrm{What\ is\ the\ nationality\ of\ Jean\mbox{-}Luc\ Godard?}$ and no directly relevant information is contained in the corpus $\mathcal{D}$, the current retrievers depend on shallow representations and lead to recall many word-level similar documents which involve $\mathrm{Jean\mbox{-}Luc\ Godard}$ or $\mathrm{Nationality}$. In this case, the retrieved information does not contribute anything semantically to answering. In fact, we know a prior the rule that if someone is born in a certain country, there is a high probability that his (her) nationality is also that country. Therefore, we can leverage this rule to conduct more effective and accurate retrieval (Huang et al., 2023) and offer documents that can better support question answering (Figure 1 (b)). Similarly, in the stage of generating answers, since we input multiple documents, LLMs may get confused by a large amount of irrelevant information. If LLMs are explicitly instructed to incorporate the retrieved content and the queries through rules, they will better attribute answers as humans expect.
Upon the above motivation, we propose Rule-Guided Retrieval-Augmented Generation (RuleRAG), a new QA approach, that enables to recall documents logically supporting queries explicitly in the directions of rules and generates the final answers based on retrieved information and attributable rules. Compared to standard RAG, training-free RuleRAG-ICL requires the introduction of rules with high confidence in the input sides of the retrievers and generators, aiming to guide the document retrieval and answer attribution processes. Moreover, to cultivate and boost the rule-following ability of LMs, we further propose RuleRAG-FT, which retrofits retrievers and generators via our designed rule-guided fine-tuning (RGFT). Specifically, for rule-guided retriever fine-tuning (RGFT-retriever), we construct pairs of queries and rules as input with the training labels of oracle retrieval examples; for rule-guided generator fine-tuning (RGFT-generator), we construct pairs of retrieved results, rules and queries as input with the training labels of golden answers. In practice, we find that our obtained documents are highly compatible with queries and the generated answers are fairly targeted with ground truths since the introduced rules improve the ability of LMs to organize causal relationships between facts (Peng et al., 2024). In summary, both RuleRAG-ICL and RuleRAG-FT can associate the retrieval stage with the generation stage, achieving better recall of information and better answer accuracy.
To demonstrate the effectiveness of RuleRAG, we newly construct five rule-aware QA benchmarks, where the queries require answering temporal or static real-world factual questions and the answers are the concrete entity names (Figure 1 is an example). It has been prone that the answers are hard to directly predict based merely on queries due to limited storage capacity (Petroni et al., 2019; Dhingra et al., 2022) and factual hallucinations (Zhang et al., 2023) of LLMs. Therefore, each benchmark offers a fact corpus which serves as the external data pool and contain documents possibly related to queries. Our experiments show that, under several retrieval and generation configurations, RuleRAG-ICL offers considerable performance gains with the individual guidance of rules by in-context learning and RuleRAG-FT achieves further improvements by combining the fine-tuned retrievers and generators. Extensive comparative studies and analyses confirm the superiority of symbolic rules and show the effectiveness of RuleRAG across various types of retrievers and generators. Moreover, RuleRAG-FT can be extrapolated to unseen rules without retraining owing to the transferable rule utilizing abilities during retrieving and generating enhanced by our designed rule-guided fine-tuning (RGFT).
Table 1: The statistics of the constructed five rule-aware QA benchmarks in this paper. $\left|\mathcal{R}\right|$, $\left|\mathcal{D}\right|$, $\left|\mathcal{F}_{R}\right|$, $\left|\mathcal{F}_{G}\right|$ and $\left|\rm{Q}\right|$ represent the numbers of rules in rule banks, documents in corpus, retriever fine-tuning query-documents pairs, generator fine-tuning query-answer pairs and test queries, respectively.
Benchmarks $|\left|\mathcal{R}\right|$$\left|\mathcal{D}\right|$ $\left|\mathcal{F}_{R}\right|$ $\left|\mathcal{F}_{G}\right|$ $\left|\rm{Q}\right|$ Query Type
RuleQA-I 557 77,508 6,594 7,440 1,559 Temporal
RuleQA-Y 99 243,633 28,153 22,765 1,864 Temporal
RuleQA-W 78 584,364 50,996 62,375 2,065 Temporal
RuleQA-F 367 49,088 8,082 9,645 1,233 Static
RuleQA-N 234 18,177 4,351 4,764 815 Static
2 Newly Constructed Rule-aware QA Benchmarks
Rule bank $\mathcal{R}$. A huge amount of world knowledge, including static facts and temporal events, has been stored in static KGs and temporal KGs (Jiang et al., 2023b). In the static scenario, several different relations can be simultaneously established between two entities. In the temporal scenario, two entities can interact multiple times at different timestamps. Hence, if relation $r_{1}$ (rule body) can logically explain the occurrence of relation $r_{2}$ (rule head) between entities, we represent this relevance as rule $r$ in a natural language form: [Entity 1, $r_{1}$, Entity 2] leads to [Entity 1, $r_{2}$, Entity 2]. We leverage the classical rule mining algorithm AMIE3 (Lajus et al., 2020) for static KGs and TLogic (Liu et al., 2022) for temporal KGs. The frequently co-occur relations form rules with high confidence (Liao et al., 2024) and we transform them to the above text string form. All these individual rules comprise our rule bank $\mathcal{R}$, which will be consistently leveraged in the training and inferring process of RuleRAG.
The statistics of our newly constructed QA benchmarks are in Table 1. Benchmarks with temporal queries, named RuleQA-I, RuleQA-Y and RuleQA-W, are constructed based on three temporal KGs, ICEWS14 (García-Durán et al., 2018), YAGO (Mahdisoltani et al., 2013) and WIKI (Leblay & Chekol, 2018). Benchmarks with static queries, named RuleQA-F and RuleQA-N, are constructed based on two static KGs, FB15K-237 (Toutanova & Chen, 2015) and NELL-995 (Xiong et al., 2017).
@@ 3 Proposed Method: RuleRAG
In this section, we present details of our proposed novel rule-guided retrieval-augmented generation with LMs (RuleRAG) for solving the task of knowledge-intensive factual queries. Notably, RuleRAG includes training-free RuleRAG-ICL and fine-tuned RuleRAG-FT. First, we prompt RuleRAG-ICL with queries and rules for in-context learning during retrieving and inferring. The rules are aimed to guide retrievers to recall logically supportive documents and guide generators to predict attributable answers. Then, RuleRAG-FT further fine-tunes the retrievers and generators to explicitly enhance their rule-following ability by our introduced rule-guided fine-tuning (RGFT), where we leverage the queries combining rules as fine-tuning data and the ground truth answers as supervision data. The inferring process of RuleRAG-FT is the same as RuleRAG-ICL.
@@ 3.1 RuleRAG-ICL
Figure 2 (a) illustrates the inference flow of RuleRAG-ICL. Given a query $q$in $\rm{Q}$, we select a few related rules $\rm{R}_{q}$ from $\mathcal{R}$. Specifically, we first recognize the relation in the query $q$ and then retrieve the rules with this relation as the rule head, forming the guidance rules $\rm{R}_{q}$ for $q$. We append $q$ with one rule $r\in\rm{R}_{q}$ once at a time to avoid conflict and conduct rule-guided retrieval in the corpus $\mathcal{D}$ to obtain the top-k relevant documents $\mathcal{D}^{r}_{q}$, where $q$ provide the retrieval content and $r$ provide retrieval directions. Finally, $\mathcal{D}^{r}_{q}$ from all rules in $\rm{R}_{q}$ are assembled to produce the final retrieval results $\mathcal{D}_{q}$, and RuleRAG-ICL conditions on the query $q$, rules $\rm{R}_{q}$ and documents $\mathcal{D}_{q}$ to reason the answer $a$.
Figure 2: The framework of our proposed RuleRAG, including RuleRAG-ICL and RuleRAG-FT. RuleRAG-ICL relies on in-context learning with the guidance of rules. RuleRAG-FT involves fine-tuning retrievers and generators ahead. (a) The unified inference process of RuleRAG. (b) Rule-guided retriever fine-tuning (RGFT-retriever). (c) Rule-guided generator fine-tuning (RGFT-generator).
Rule-guided retriever (RG-retriever). Since each rule $r\in\rm{R}_{q}$ stands for a unique retrieval logic, we retrieve all query-rule pairs ($q$,$r$) individually to avoid the conflict of different rules. Specifically, the retriever calculates a relevant score $s(d_{i},q\circ r)$ between a ($q$,$r$) pair and every document $d_{i}\in\mathcal{D}$: $s(d_{i},q\circ r)=\mathbf{E}_{d}(d_{i})\cdot\mathbf{E}_{q}(q\circ r)$, where $\circ$ denotes sequence concatenation, $\cdot$ is dot product, $\mathbf{E}_{d}$ is the document encoder and $\mathbf{E}_{q}$ is the query encoder. To stay within the context window size limit of LLMs, we select the top-k scored documents, denoted as $\mathcal{D}_{q}^{r}$ ($r\in\rm{R}_{q}$), for each ($q$,$r$) pair and combine all $\mathcal{D}_{q}^{r}$ as the final retrieval results $\mathcal{D}_{q}$ for query $q$. This process is formalized as follows:
$\mathcal{D}_{q}=\textstyle\bigcup\limits_{r\in\rm{R}_{q}}\mathcal{D}_{q}^{r},\ \left(\rm{R}_{q}\subsetneqq\mathcal{R}\right);\mathcal{D}_{q}^{r}=\text{arg}\ \ {\underset{d_{i}\in\mathcal{D}}{\text{top}\mbox{-}\text{k}}}\ \ s\left(d_{i},q\circ r\right),\ (r\in\rm{R}_{q}).$ (1)
Rule-guided generator (RG-generator). After recalling $\mathcal{D}_{q}$, we construct an instruction to prompt LLMs to generate the final answer $a$. Different from the widely used case-based prompts (Wei et al., 2024), we do not let LLMs learn the reasoning mode implicitly from examples, but directly inform LLMs of $\rm{R}_{q}$ as the attribution mechanisms and make LLMs answer the query $q$ explicitly according to the $\mathcal{D}_{q}$. Under the guidance of rules, the probability of outputting $a$ can be approximated as follows:
$P(a\mid q)=P(a\mid q,\mathcal{R},\mathcal{D})\approx P_{LLM}(a\mid\textsc{instruction}(q,\rm{R}_{q},\mathcal{D}_{q})),$ (2)
where $P_{LLM}\textsc{()}$) is the generation probability of LLMs and instruction() is the instruction prompt. The simplified form of the instruction is in Figure 2 (c) and the detail is given in the Appendix A.9.

### v2

@@ Abstract
Retrieval-augmented generation (RAG) footnotetext: * Corresponding Authors has shown promising potential in knowledge intensive question answering (QA). However, existing approaches only consider the query itself, neither specifying the retrieval preferences for the retrievers nor informing the generators of how to refer to the retrieved documents for the answers, which poses a significant challenge to the QA performance. To address these issues, we propose Rule-guided Retrieval-Augmented Generation with LMs, which explicitly introduces rules for in-context learning (RuleRAG-ICL) to guide retrievers to recall related documents in the directions of rules and uniformly guide generators to reason attributed by the same rules. The combination of queries and rules can be used as fine-tuning data to update retrievers and generators, achieving better rule-based instruction-following ability (RuleRAG-FT). Moreover, most existing RAG datasets were constructed without considering rules and Knowledge Graphs (KGs) are recognized as providing high-quality rules. Therefore, we construct five rule-aware RAG benchmarks for QA, RuleQA, based on KGs to stress the significance of retrieval and reasoning with rules. Experiments on RuleQA demonstrate RuleRAG-ICL improves the retrieval quality of +89.2% in Recall@10 and answer accuracy of +103.1% in Exact Match, and RuleRAG-FT yields more enhancement. In addition, experiments on four existing RAG datasets show RuleRAG is also effective by offering rules in RuleQA to them, further proving the generalization of rule guidance in RuleRAG. Code and RuleQA are at https://anonymous.4open.science/r/RuleRAG.
RuleRAG: Rule-Guided Retrieval-Augmented Generation with Language Models for Question Answering
Zhongwu Chen1, Chengjin Xu2∗, Dingmin Wang3∗, Zhen Huang1∗, Yong Dou1, Jian Guo2 1National Key Laboratory of Parallel and Distributed Computing, 1College of Computer Science and Technology, National University of Defense Technology 2International Digital Economy Academy (IDEA) 3Department of Computer Science, University of Oxford {chenzhongwu20, huangzhen, yongdou}@nudt.edu.cn, {xuchengjin, guojian}@idea.edu.cn, {dingmin.wang}@cs.ox.ac.uk
@@ 1 Introduction
Large language models (LLMs) have achieved the impressive capability of language generation and knowledge learning (Brown et al., 2020; Ouyang et al., 2022). Despite the success, the full-parametric knowledge in LLMs struggles to precisely manipulate fine-grained queries, especially in knowledge-intensive tasks (Jiang et al., 2023c; Shao et al., 2023). As complementary, RAG shows superior performance in many NLP tasks, such as open-domain QA (Trivedi et al., 2023) and natural language inference (Qin et al., 2023).
@@
Figure 1: (a) Without the help of rules, the current RAG can only retrieve documents about some keywords, rather than the overall semantics of the query, and thus get confused in answering. (b) Guided by the attributable rule $r$, our proposed RuleRAG retrieves logically supportive documents and then reasons the correct answer .
@@
However, two high-level issues exist in the current RAG. First, in the retrieval phase, the retrievers rely on word-level matching, and thus can not guarantee that the recalled information is always pertinent to the query answering. The reason is many retrievers are trained on unsupervised text or trained end-to-end, leading to their insufficiency in retrieving the necessary statements for reasoning (BehnamGhader et al., 2023). Secondly, in the generation phase, the LLMs in the current RAG are not specifically informed of how to exploit noisy retrieved content properly, since relationships between a wide range of facts are rarely explicitly “pointed out” and “supervised” in the pre-training corpora of LLMs. Even if answered correctly, they still lead to implicit attribution processes that are difficult to explain and verify. Therefore, the current RAG is neither inherently trained to retrieve along reasonable retrieval directions nor organically attribute retrieved content to answers.
While answering knowledge-intensive queries, a priori rules instead of simply matching words can genuinely capture internal logical patterns among complicated knowledge. Some works incorporate rules into LLMs to handle the addition of numbers (Hu et al., 2024) or industrial tasks (Zhang et al., 2024b), but there is currently no exploration of introducing rules into RAG for QA. As shown in Figure 1, the query is What is the trend in YSSTECH’s stock price going forward?. Current retrievers recall many documents that contribute nothing to answering because of blind retrieval. By contrast, financial KGs provide a rule that The merger of a company’s businesses with other influential companies leads to The increase in a company’s stock price. Therefore, we can leverage this rule to conduct more targeted retrieval and offer documents that can better support question answering.
Upon the above motivation, we propose RuleRAG, Rule-guided Retrieval-Augmented Generation, which enables to both retrieve documents and reason answers with the guidance of rules. Compared to standard RAG, which relies on finding the precise statement of the query to be answered, training-free RuleRAG-ICL requires the introduction of rules in the input sides of the retrievers to retrieve and generators to reason. To boost the rule-following reasoning ability, we further design rule-guided fine-tuning (RGFT) to retrofit retrievers and generators (RuleRAG-FT).
However, although rules are common and valuable for the QA task, most existing RAG datasets were constructed without considering rules. KGs are widely known to provide question-answer pairs and high-coverage rules, so we newly construct five rule-aware QA benchmarks (RuleQA) from five KGs to offer rich data. The strong performance of KG rule mining algorithms also ensures the acquisition of rules with high confidence. In addition, the answers in RuleQA require reasoning based on rules rather than directly repeating retrieved facts like existing RAG datasets, so RuleQA is knowledge-intensive and more challenging. Experiments on RuleQA show that, under several retrieval and generation configurations, RuleRAG-ICL offers considerable performance gains with the individual guidance of rules by in-context learning and RuleRAG-FT achieves further improvements by fine-tuning retrievers and generators. RuleRAG-FT can also be extrapolated to unseen rules without retraining, confirming the superiority of rules to retrieve and reason in RAG. Moreover, we conduct experiments on four existing RAG datasets by adaptively equipping our mined rules for them. We find both RuleRAG-ICL and RuleRAG-FT can improve their performance, implying that apart from our constructed benchmarks RuleQA, it is also beneficial to introduce rules to existing RAG datasets.
2 Related Work
2.1 Retrieval-Augmented Generation
In RAG, the retrieval module explicitly augments the generation module with external knowledge banks (Guu et al., 2020). Retrieval approaches include sparse retrievers based on sparse bag-of-words, dense retrievers based on dense vectors and more complex hybrid search algorithms (Li et al., 2023a). RAG is widely adopted to complement the LLM parametric knowledge along different stages (Gao et al., 2024), including pre-training stage (RETRO (Borgeaud et al., 2022), COG (Lan et al., 2023), Atlas (Izacard et al., 2024)), fine-tuning stage (SURGE (Kang et al., 2023), Self-RAG (Asai et al., 2023), CoN (Yu et al., 2023)) and inference stage (KnowledGPT (Wang et al., 2023), DSP (Khattab et al., 2023), RoG (Luo et al., 2024)).
Figure 2: The framework of our proposed RuleRAG. RuleRAG-ICL relies on in-context learning with the guidance of rules. RuleRAG-FT involves fine-tuning retrievers and generators ahead. (a) The unified RuleRAG inference process. (b) Rule-guided retriever fine-tuning (RGFT-retriever). (c) Rule-guided generator fine-tuning (RGFT-generator).
2.2 Knowledge-intensive QA
In the realm of QA, queries are viewed as knowledge-intensive if we need access to external corpora (Thorne et al., 2018). Assuming that documents in the corpora include the exact answers, RAFT (Zhang et al., 2024a) and RA-DIT (Lin et al., 2024) fine-tune LLMs by concatenating documents and queries as prompts. However, many answers to factual queries are hidden in semantically dissimilar but logically related documents, which need to be retrieved and reasoned with the guidance of rules. Our constructed RuleQA simulates this circumstance, while most existing RAG datasets lack rules. Recently,  Wu et al. (2024) investigates mitigating misleading irrelevant interference;  Sun et al. (2024) only discusses the rule-following abilities of LLMs without retrieval and ignores how to obtain rules. In contrast, our proposed RuleRAG involves a more comprehensive consideration of mining rules, retrieving documents and reasoning answers.
@@ 3.1 RuleRAG-ICL
Figure 2 (a) illustrates the inference flow of RuleRAG-ICL. Given a query $q$in $\rm{Q}$, we first leverage Sentence-BERT (Reimers and Gurevych, 2019) to capture the semantic similarity between $q$ and candidate rules. The highest N rules among those whose scores exceed a certain threshold $\theta$ are taken as guiding rules $\rm{R}_{q}$, where N and $\theta$ are hyper-parameters. Then, we append $q$ with one rule $r\in\rm{R}_{q}$ once at a time to avoid conflict and conduct rule-guided retrieval in the corpus $\mathcal{D}$ to obtain the top-k documents $\mathcal{D}^{r}_{q}$. Finally, $\mathcal{D}^{r}_{q}$ from all rules in $\rm{R}_{q}$ are assembled to produce the final retrieval results $\mathcal{D}_{q}$, and RuleRAG-ICL conditions on the query $q$, rules $\rm{R}_{q}$ and documents $\mathcal{D}_{q}$ to reason the answer $a$.
Rule-guided retriever (RG-retriever). The retriever calculates a relevant score $s(d_{i},q\circ r)$ between ($q$,$r$) and every document $d_{i}\in\mathcal{D}$: $s(d_{i},q\circ r)=\mathbf{E}_{d}(d_{i})\cdot\mathbf{E}_{q}\left(q\circ r\right)$, where $\circ$ is sequence concatenation, $\cdot$ is dot product, $\mathbf{E}_{d}$ is document encoder and $\mathbf{E}_{q}$ is query encoder. We select the top-k documents, $\mathcal{D}_{q}^{r}$, and combine all $\mathcal{D}_{q}^{r}$ as the final retrieval results $\mathcal{D}_{q}$. This process is formalized as:
$\mathcal{D}_{q}=\textstyle\bigcup\limits_{r\in\rm{R}_{q}}\mathcal{D}_{q}^{r};$ $\mathcal{D}_{q}^{r}=\text{arg}\ \ {\underset{d_{i}\in\mathcal{D}}{\text{top}\mbox{-}\text{k}}}\ \ s\left(d_{i},q\circ r\right).$ (1)
Rule-guided generator (RG-generator). After recalling $\mathcal{D}_{q}$, we construct an instruction to prompt LLMs to reason answers. Different from implicit case-based prompts (Wei et al., 2024), we directly inform LLMs of $\rm{R}_{q}$ as the attribution mechanisms and reason with the guidance of $\rm{R}_{q}$. The probability of outputting answer $a$ can be approximated as:
$P(a\mid q,\mathcal{R},\mathcal{D})\approx P_{LLM}(a\mid\textsc{ins}(q,\rm{R}_{q},\mathcal{D}_{q})),$ (2)
where $P_{LLM}\textsc{()}$) is the LLM generation probability. ins() is instruction prompt, whose simplified form is in Figure 2 (c) and details are in Appendix K.